OpenStudy (anonymous):

two forces, with magnitudes 13N and 9N, act on an object. the magnitude of the resultant force is 19.16N. find the angle at which the two forces are acting on each other

OpenStudy (asnaseer):

 use:\[\tan(\alpha)=\frac{13}{9}\]and\[\tan(\beta)=\frac{9}{13}\]to find the angles \(\alpha\) and \(\beta\).